OpenStudy (anonymous):

I need some help! (WILL GIVE MEDALS) what is x in 10x=2

rvc (rvc):

1/5 x=2/10 1/5

OpenStudy (anonymous):

Start by dividing each term in the equality by 10

OpenStudy (anonymous):

I do that one sec.

OpenStudy (anonymous):

This is the division property of equality right?

rvc (rvc):

orelse send 10 to otherside which becomes division

OpenStudy (anonymous):

It's really simple. 10x/10=2/10 and yes

OpenStudy (anonymous):

What happens to the 10s on the right side?

OpenStudy (anonymous):

x equals .2

OpenStudy (anonymous):

No, the 10s cancel each other out and you are left with x=2/10 which you then simplify to get what?

OpenStudy (anonymous):

10x=2 x equals .2 10 times .2 equals 2 THANK YOU GUYS SO MUCH.

OpenStudy (anonymous):

You're welcome but just in case, make x=1/5 because that is likely the answer your teacher wants

rvc (rvc):

did u get it will u do 1 more @Pcx16

OpenStudy (anonymous):

OK

rvc (rvc):

right 24x=8

OpenStudy (anonymous):

x= 5/10 that equals .2

OpenStudy (anonymous):

rvc let me try that one give me a sec.

OpenStudy (ipwnbunnies):

5/10 is equal to .5. 5/25 equals .2

rvc (rvc):

wat?

OpenStudy (ipwnbunnies):

I just wanted to correct PCX. He said 5/10 equals 0.2

OpenStudy (anonymous):

x equals .33 rvc thats then 24x equals 8 24 times .33 equals 8

OpenStudy (anonymous):

Ipwn that was just a problem that britts wanted me to do. THANK YOU ALL FOR THE HELP!

rvc (rvc):

its wrong @Pcx16

rvc (rvc):

right @Pcx16 sorry didn't see the point